Photoelectric threshold for a metal is 300 nm. Determine the maximum energy of electrons emitted by a radiation of 200 nm
1
Expert's answer
2020-09-29T09:40:27-0400

We can find the maximum kinetic energy of electrons from the formula:


"KE_{max} = h \nu - W = \dfrac{hc}{\lambda} - W,"

here, "KE_{max}" is the maximum kinetic energy of electrons, "h = 6.63 \cdot 10^{-34} \ J \cdot s" is the Planck's constant, "\nu = \dfrac{c}{\lambda}" is the frequency of the electromagnetic radiation, "c = 3 \cdot 10^8 \ \dfrac{m}{s}" is the speed of light, "\lambda = 200 \ nm" is the length of the electromagnetic radiation and "W" is the work function.

Let's find the work function:


"W= h \nu_0 = \dfrac{hc}{\lambda_0},"


here, "\lambda_0 = 300 \ nm" is the threshold wavelength.

Then, we get:


"W= \dfrac{6.63 \cdot 10^{-34} \ J \cdot s \cdot 3 \cdot 10^8 \ \dfrac{m}{s}}{300 \cdot 10^{-9} \ m} = 6.63 \cdot 10^{-19} \ J."

Finally, we can find the maximum kinetic energy of electrons:

"KE_{max} = \dfrac{6.63 \cdot 10^{-34} \ J \cdot s \cdot 3 \cdot 10^8 \ \dfrac{m}{s}}{200 \cdot 10^{-9} \ m} - 6.63 \cdot 10^{-19} \ J = 3.315 \cdot 10^{-19} \ J."

Answer:

"KE_{max} = 3.315 \cdot 10^{-19} \ J."